OpenStudy (anonymous):

The area of a triangle is 67. If the height was 13, find the value of the base.

OpenStudy (anonymous):

20.62 9.89 10.31 2.58 ?

OpenStudy (anonymous):

area \(= \frac 12 Base \space \times \space Height \)

OpenStudy (anonymous):

can you do it from here ?

OpenStudy (nubeer):

just use Area=1/2x(height)x(base) you will get it just try yourself

OpenStudy (anonymous):

yes, basically what i have to do know is solve x*13=67

OpenStudy (anonymous):

don't forget the 1/2

OpenStudy (anonymous):

10.31

OpenStudy (anonymous):

oh yeah! I didn't see that

OpenStudy (anonymous):

area of a triangle is (1/2)(base)(height).

OpenStudy (anonymous):

I got 2.58

OpenStudy (anonymous):

nope, can you show your steps?

OpenStudy (anonymous):

oh no, i just found out that is incorrect and I already submitted it. Do you still want to know what I did?

OpenStudy (anonymous):

only if you want to learn :-)

OpenStudy (anonymous):

1/2bh=67 b=67*2/h b=134/13 b=10.31

OpenStudy (anonymous):

I did 67/13=5.15 5.15/.5= ~2.58

OpenStudy (anonymous):

when you divide by 0.5 that's multiplying by 2. you divided by 2.

OpenStudy (anonymous):

I need to multiply didn't I?

OpenStudy (anonymous):

*needed

OpenStudy (anonymous):

abhed showed the right way

OpenStudy (anonymous):

you wrote "5.15 / .5 = 2.58", that's not correct

OpenStudy (anonymous):

Yeah, thanks abhed!

OpenStudy (anonymous):

5.15 / .5 = 10.3

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

yeah I see... well at least I won't make the same mistake again... i hope. Thanks you guys! Really, you've helped a lot.